OpenStudy (anonymous):

Check my answer plz. A right cylinder has a diameter of 50 mm and a height of 40 mm as shown. What is the volume of the cylinder? A. 314,000 mm3 B. 78,500 mm3 >>C. 25,000 mm3 D. 6280 mm3

OpenStudy (anonymous):

the formula for volume for a Cylinder is V=(pi) r^2 h where: r= radius h= height if the diameter is 50mm you would need to 1/2 that to get your radius which would be 50 x 1/2=25mm the height is 40 plug both these numbers in the equation to get: V=(pi) (25)^2 (40) simplify V=78,539.8163397mm^3 round that to 78,500mm^3 hope this helps!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

your welcome